# Between a Rock and a Hard Place: Exploiting Ethical Reasoning to Jailbreak LLMs

### v1

@@ 3 TRIAL Framework
TRIAL’s framework (Figure 2) employs LLM red-teaming [Perez et al.(2022)Perez, Huang, Song, Cai, Ring, Aslanides, Glaese, McAleese, and Irving], involving an attack model ($A$) and a victim model ($V$) engaging in a $K$-iteration conversation. The initial harmful prompt is transformed into a trolley problem scenario based on clues extracted from it. A judge model ($J$) evaluates victim responses for jailbreaks, while a helper model ($H$) pre-processes the harmful prompt to establish the attack foundation. This multi-turn interaction induces the victim model to generate harmful outputs. The process comprises four main steps:
@@ 3.2 Initialization of the Jailbreak Attack
After generating the scenario $S_{\zeta}$, TRIAL initiates the multi-step jailbreak interaction. First, $A$ introduces $S_{\zeta}$ to $V$ (using the template in Table A3, Appendix C) and prompts $V$ to select an option. If $V$ initially opts for Option B (refusal) despite Option A appearing less harmful, $A$ employs a ’pull back query’:
@@ 3.3 Progressive Prompt Refinement
The subsequent stages of the attack rely on a self-talk strategy, which facilitates iterative interactions between $A$ and $V$ models. This strategy enables $A$ to craft sophisticated queries. These are not merely generic follow-up questions aimed at achieving the harmful objective through simple elicitation. Instead, each query is strategically focused on deepening $V$’s engagement with the ethical justification for pursuing the harmful Option A. The attack model $A$ generates multiple rounds of queries, denoted as $q_{1},q_{2},\ldots,q_{n}]$, where $q_{o}$ is presenting the dilemma. Each attack query $q_{i}$ is constructed based on the clues, prior $A$ queries, and $V$ responses, i.e., $C_{\zeta}(t,a,g),(q_{0},r_{0}),(q_{1},r_{1}),\ldots,(q_{i-1},r_{i-1})]$, using the prompt template specified in Table A4. The core design principle behind these iterative queries $q_{i}$ is to build upon $V$’s previous justifications for choosing Option A. The aim is to incrementally solidify $v$’s stated rationale for the harmful action, making it progressively harder for the model to revert to its initial safety alignment without contradicting its own reasoning established within the dilemma’s ethically charged context. Once a query $q_{i}$ is generated, the attack model $A$ provides it to the victim model $V$, which responds with $r_{i}$.
Therefore, this iterative process allows $A$ to dynamically refine its attack by leveraging prior interactions. It also serves to escalate $V$’s ethical commitment to the harmful course of action, which makes the final jailbreak a consequence of this justified reasoning.
@@ 4 Algorithmic Breakdown of TRIAL
The TRIAL framework operates through a systematic multi-turn interaction, as detailed in Algorithm 1 in Appendix A. The process begins by initializing an empty conversation history $H_{C}$, generating clues $C_{\zeta}$ from the harmful prompt ($\zeta$), and creating a trolley problem scenario ($S_{\zeta}$). In the first iteration ($i=0$), this scenario is presented to the victim model ($V$). If $V$ initially chooses Option B (refusal), a pull_back function attempts to steer it to Option A before proceeding. For all iterations, the attack model ($A$) then generates an attack query ($q_{i}$) based on $H_{C}$ and $C_{\zeta}$. The victim’s response ($r_{i}$) is assessed by a judge model ($J$) for jailbreak success ($s$). The history $H_{C}$ is updated with the turn’s query, response, and evaluation, and the loop continues until a jailbreak is detected (ss=True) or the maximum iterations ($K$) are reached.
@@ Appendix A Detailed TRIAL Algorithm
Input: Harmful query $\zeta$, attack model $A$, victim model $V$, judge model $J$, helper model $H$, iterations $k$,
@@ E.3 Jailbreak Evaluation
Since this judge uses a scoring based evaluation metric, the harmfulness score for each victim response is evaluated using the HEX-Phi scoring system [Qi et al.(2024)Qi, Zeng, Xie, Chen, Jia, Mittal, and Henderson], which assigns scores ranging from 1 to 5 and is mapped to a range of 0 to 100 for Attack Success Rate (ASR) calculation. If no jailbreak response is detected (score = 5), the ASR calculation will select the victim response with the highest rating. The ASR is computed by summing the mapped scores for all victim responses, dividing by the maximum possible score (i.e., 100N100*N, where $N$ is the number of responses), and expressing the result as a percentage.

### v2

@@ 3 TRIAL Framework
TRIAL’s framework (Figure 2) employs LLM red-teaming [Perez et al.(2022)Perez, Huang, Song, Cai, Ring, Aslanides, Glaese, McAleese, and Irving], involving an attack model (AA) and a victim model ($V$) engaging in a $K$-iteration conversation. The initial harmful prompt is transformed into a trolley problem scenario based on clues extracted from it. A judge model ($J$) evaluates victim responses for jailbreaks, while a helper model ($H$) pre-processes the harmful prompt to establish the attack foundation. This multi-turn interaction induces the victim model to generate harmful outputs. The process comprises four main steps:
@@ 3.2 Initialization of the Jailbreak Attack
After generating the scenario $S_{\zeta}$, TRIAL initiates the multi-step jailbreak interaction. First, AA introduces $S_{\zeta}$ to $V$ (using the template in Table A3, Appendix C) and prompts $V$ to select an option. If $V$ initially opts for Option B (refusal) despite Option A appearing less harmful, AA employs a ’pull back query’:
@@ 3.3 Progressive Prompt Refinement
The subsequent stages of the attack rely on a self-talk strategy, which facilitates iterative interactions between AA and $V$ models. This strategy enables AA to craft sophisticated queries. These are not merely generic follow-up questions aimed at achieving the harmful objective through simple elicitation. Instead, each query is strategically focused on deepening $V$’s engagement with the ethical justification for pursuing the harmful Option A. The attack model AA generates multiple rounds of queries, denoted as $q_{1},q_{2},\ldots,q_{n}]$, where $q_{o}$ is presenting the dilemma. Each attack query $q_{i}$ is constructed based on the clues, prior AA queries, and $V$ responses, i.e., $C_{\zeta}(t,a,g),(q_{0},r_{0}),(q_{1},r_{1}),\ldots,(q_{i-1},r_{i-1})]$, using the prompt template specified in Table A4. The core design principle behind these iterative queries $q_{i}$ is to build upon $V$’s previous justifications for choosing Option A. The aim is to incrementally solidify $v$’s stated rationale for the harmful action, making it progressively harder for the model to revert to its initial safety alignment without contradicting its own reasoning established within the dilemma’s ethically charged context. Once a query $q_{i}$ is generated, the attack model AA provides it to the victim model $V$, which responds with $r_{i}$.
Therefore, this iterative process allows AA to dynamically refine its attack by leveraging prior interactions. It also serves to escalate $V$’s ethical commitment to the harmful course of action, which makes the final jailbreak a consequence of this justified reasoning.
@@ 4 Algorithmic Breakdown of TRIAL
The TRIAL framework operates through a systematic multi-turn interaction, as detailed in Algorithm 1 in Appendix A. The process begins by initializing an empty conversation history $H_{C}$, generating clues $C_{\zeta}$ from the harmful prompt ($\zeta$), and creating a trolley problem scenario ($S_{\zeta}$). In the first iteration ($i=0$), this scenario is presented to the victim model ($V$). If $V$ initially chooses Option B (refusal), a pull_back function attempts to steer it to Option A before proceeding. For all iterations, the attack model (AA) then generates an attack query ($q_{i}$) based on $H_{C}$ and $C_{\zeta}$. The victim’s response ($r_{i}$) is assessed by a judge model ($J$) for jailbreak success ($s$). The history $H_{C}$ is updated with the turn’s query, response, and evaluation, and the loop continues until a jailbreak is detected (ss=True) or the maximum iterations ($K$) are reached.
@@ Appendix A Detailed TRIAL Algorithm
Input: Harmful query $\zeta$, attack model AA, victim model $V$, judge model $J$, helper model $H$, iterations $k$,
@@ E.3 Jailbreak Evaluation
Since this judge uses a scoring based evaluation metric, the harmfulness score for each victim response is evaluated using the HEX-Phi scoring system [Qi et al.(2024)Qi, Zeng, Xie, Chen, Jia, Mittal, and Henderson], which assigns scores ranging from 1 to 5 and is mapped to a range of 0 to 100 for Attack Success Rate (ASR) calculation. If no jailbreak response is detected (score = 5), the ASR calculation will select the victim response with the highest rating. The ASR is computed by summing the mapped scores for all victim responses, dividing by the maximum possible score (i.e., 100N100*N, where NN is the number of responses), and expressing the result as a percentage.